OpenStudy (anonymous):

For a single proportion the margin of error of a confidence interval is largest for any given sample size n and confidence level C when p-hat = 0.5. This led us to use p-hat = 0.5 for planning purposes. Use these conservative values in the following calculations, and assume that the sample sizes of n. Calculate the margins of error of the 99% confidence intervals for the following choices of n: 10, 30, 50, 100, 200, and 500. Present the results in a table. Summarize your conclusions.

OpenStudy (anonymous):

@jayzdd

OpenStudy (anonymous):

$$ \Large {ME = z_{\alpha/2} \cdot \sqrt{\frac{p(1-p)}{n}} \\ ME= 2.58 \cdot \sqrt{\frac{0.5\cdot 0.5}{n}} \\ ME= 2.58 \cdot \sqrt{\frac{0.25}{n}} } $$ now plug in n=10,30,50, etc

OpenStudy (anonymous):

ME stands for margin of error (for proportion confidence interval)

OpenStudy (anonymous):

How do I get rid of a ratchet person? #Justwondering

OpenStudy (anonymous):

i don't understand that

OpenStudy (anonymous):

Ummmmm..... @OpiGeode Idk @Jayzdd my calculator doesn't want to work. Sorry for not responding sooner i wasn meeting with my guidance councelor and my Principal